### v1

@@ Abstract
Long time-duration low-thrust nonlinear optimal spacecraft trajectory global search is a computationally and time expensive problem characterized by clustering patterns in locally optimal solutions. During preliminary mission design, mission parameters are subject to frequent changes, necessitating that trajectory designers efficiently generate high-quality control solutions for these new scenarios. Generative machine learning models can be trained to learn how the solution structure varies with respect to a conditional parameter, thereby accelerating the global search for missions with updated parameters. In this work, state-of-the-art diffusion models are integrated with the indirect approach for trajectory optimization within a global search framework. This framework is tested on two low-thrust transfers of different complexity in the circular restricted three-body problem. By generating and analyzing a training data set, we develop mathematical relations and techniques to understand the complex structures in the costate domain of locally optimal solutions for these problems. A diffusion model is trained on this data and successfully accelerates the global search for both problems. The model predicts how the costate solution structure changes, based on the maximum spacecraft thrust magnitude. Warm-starting a numerical solver with diffusion model samples for the costates at the initial time increases the number of solutions generated per minute for problems with unseen thrust magnitudes by one to two orders of magnitude in comparison to samples from a uniform distribution and from an adjoint control transformation.
@@ 1 Introduction
Low-thrust trajectory optimization is a global search problem with multiple objectives and constraints. Finding optimal solutions with respect to the primary objectives of minimizing fuel consumption and time of flight is a complex, computationally expensive and time-consuming endeavour tackled in the preliminary mission design phase. This task is further complicated by frequent changes in mission parameters and objectives. The flight dynamics team must often adapt quickly to such changes and generate new optimal solutions. For example, adjustments to the maximum thrust magnitudes, typically reported by the propulsion team, can lead to significantly different optimal trajectories. Such adjustments might also be required mid-mission, as exemplified by the BepiColombo spacecraft, operated by ESA and JAXA. In May 2024, a thruster glitch limited its operation to $90\%$ of maximum power [1]. Ideally, control solutions for such new constraints would be generated swiftly, either by ground engineers or autonomous on-board algorithms. We present a framework that uses generative machine learning to speed up the global search for trajectory optimization problems with varying mission parameters.
@@
Previous studies have demonstrated that locally optimal solutions to low-thrust trajectory optimization problems tend to form clusters [2, 3, 4, 5, 6]. Global search algorithms like monotonic basin hopping (MBH) aim to exploit this clustering to improve efficiency in generating solutions [7]. However, MBH requires manually tuning the sampling distributions, which are typically chosen for their simplicity such as uniform, Gaussian, Cauchy, or Pareto distributions. Li et al. [2] proposed an amortized global search framework that learns a conditional probability distribution supported on local basins of attraction for high-quality solutions using deep generative models. Their distribution was conditioned on the spacecraft’s maximum thrust, initially employing a conditional variational autoencoder (CVAE), later enhanced with a long short-term memory (LSTM), alongside a numerical solver to generate solutions for new thrust values. Beeson et al. [3] built on this work by mathematically defining the parameterized global search problem as the task of sampling a conditional probability distribution, which focuses on basins of attractions of high-quality solutions. Subsequent studies [8] found that diffusion models (DM) outperformed the CVAE + LSTM approach for trajectory optimization. Diffusion probabilistic models, are state-of-the-art generative models, that have recently captured widespread interest, due to there strong performance in conditional image and video generation [9].
@@
The novel contribution of this work is the integration of a diffusion model framework with indirect optimal control for spacecraft trajectory optimization. While this builds on the work of Li et al. [2, 8], their research has focused solely on direct optimal control methods. Direct methods use transcription techniques to approximate dynamics and control, reformulating the optimal control problem as a nonlinear program. In contrast, indirect methods derive the first-order necessary conditions for optimality, where the additional costate variables are introduced. This results in a two-point boundary value problem. Although direct methods are now the standard in spacecraft trajectory optimization due to their simpler setup and robust convergence properties, especially with poor initial guesses, indirect methods provide a lower-dimensional solution space where each solution satisfies the first-order optimality conditions [10]. The main difficulty with the indirect method lies in generating good initial guesses for the non-intuitive costate variables, which are crucial for solver convergence. This challenge drives the use of generative learning in this context. By training a diffusion model to learn the structure of high-quality solutions in costate space, it can generate initial costate guesses for new conditional parameters.
@@
While the diffusion model facilitates the rapid generation of solutions for new conditional parameters, it requires a substantial amount of data for training. A key contribution of this work is the implementation of efficient methods to generate this training data. This is achieved by integrating the astrodynamics software pydylan [15] with adjoint control transformations (ACTs) and a modified version of the preliminary screening algorithm introduced by Russell [16]. The generative learning framework’s capabilities are demonstrated through two low-thrust transfers within the circular restricted three-body problem (CR3BP), with both transfers conditioned on the spacecraft’s maximum thrust. Training data is generated for fixed values of this parameter, and the diffusion model is then tested by generating solutions for a new thrust level not included in the training set. Results show that the diffusion model accelerates the global search process by one to two orders of magnitude, allowing for rapid generation of high-quality solutions for the new problem.
@@ 2.1 Optimal Control Problem
The goal of optimal control theory is to determine a control signal which causes a process to satisfy a set of physical constraints, while maximizing or minimizing a performance measure [17]. This performance index is formulated as a cost or objective functional $J$, including a terminal cost $\phi$ and a running cost $\mathcal{L}$. A generic formulation of the problem is given by:
@@ 2.4 Circular Restricted Three Body Problem
Selecting a suitable dynamical model is vital to space mission design, as each strikes a different balance between realism and complexity. The circular restricted three-body problem (CR3BP) is commonly used during preliminary mission design for multi-body missions, as it simplifies the problem while utilizing dynamical structures for effective mission planning. In the CR3BP, a primary celestial body of mass $m_{1}$ and a secondary celestial body of mass $m_{2}$ ($m_{1}>m_{2}$) move in circular orbits around their common center of mass. The CR3BP describes the motion of a third body of negligible mass, that does not affect the orbits of the larger bodies [21]. Variables in the CR3BP are expressed using a system of normalized units, referred to as natural units (NU): the unit of mass is $m_{1}+m_{2}$, the distance unit (DU) is the constant distance between primary and secondary, and the time unit (TU) is chosen such that their orbital period around the barycenter is $2\pi$. The system’s only parameter is the mass ratio $\mu=\frac{m_{2}}{m_{1}+m_{2}}$. The third body’s motion is described in a rotating reference frame, where the primary and secondary are fixed along the $r_{1}$axis at $(-\mu,0,0)$ and $(1-\mu,0,0)$, with their barycenter at the origin. The $r_{3}$axis points in the direction of angular momentum, while the $r_{2}$-axis completes the right-handed system. The motion of the third body under the gravitational influence of $m_{1}$ and $m_{2}$ follows the dynamical system:
@@ 3.1 Adjoint Control Transformation
When using an indirect method, an initial guess for the costate variables is required to solve the boundary value problem. This is challenging due to the small convergence radius and lack of physical intuition for costates [23]. For data generation we employ a method referred to as adjoint control transformation (ACT) [24], which maps a set of physical control variables to the non-physical costates: $\mathcal{M}:(\varphi,\dot{\varphi},\beta,\dot{\beta},S,\dot{S})\rightarrow(\boldsymbol{\lambda}_{r},\boldsymbol{\lambda}_{v})$, where $\varphi$ and $\beta$ represent the in-plane and out-of-plane angles of the thrust vector in a velocity-centered spacecraft coordinate frame. We evaluate this mapping at $t_{0}$ by uniformly sampling initial guesses for $\varphi_{0}$, $\beta_{0}$, $\dot{\varphi}_{0}$, $\dot{\beta}_{0}$, $S_{0}$ and $\dot{S}_{0}$ within problem-specific ranges, yielding reasonable initial guesses for $\boldsymbol{\lambda}_{r,0}$ and $\boldsymbol{\lambda}_{v,0}$. No mapping is needed for the mass costate at $t_{0}$, as we exploit the degree of freedom introduced by the factor $k$ by setting $\lambda_{m,0}=-1$. The transversality condition $\lambda_{m}(t_{f})=-k$ is satisfied, since $\dot{\lambda}_{m}\leq 0$.
$\hat{\boldsymbol{v}}$$\hat{\boldsymbol{w}}$$\hat{\boldsymbol{h}}$$\boldsymbol{v}$$\boldsymbol{u}$$\beta$$\varphi$$r_{1}$$r_{2}$$r_{3}$$m_{1}$$m_{2}$
@@ 3.2 Global Search Strategy
1. Set fixed values: $\boldsymbol{r}(0)=\boldsymbol{r}_{0}$, $\boldsymbol{v}(0)=\boldsymbol{v}_{0}$, $m(0)=m_{0}$, $\lambda_{m,0}=-1$, $\tau_{i}=0$ 2. Generate k-d tree for states on target orbit or manifold 3. DO WHILE ($j\leq j_{max}$ with $j\in\mathbb{N}$) (a) $j=j+1$ (b) Apply ACT method to obtain $\boldsymbol{\lambda}_{r,0}$ and $\boldsymbol{\lambda}_{v,0}$ (c) Propagate spacecraft states and costates according to Eqs. (6), (11) and (12) for $\tau_{s,max}$ and save $\boldsymbol{r}(t)$ and $\boldsymbol{v}(t)$ (d) $\delta_{c,min}=\infty$ (e) FOR $t=0$ to $t=\tau_{s,max}$ DO i. Query k-d tree with $\begin{pmatrix}\boldsymbol{r}(t)\\ \boldsymbol{v}(t)\end{pmatrix}$ ii. k-d tree returns closest point on target orbit or manifold $\begin{pmatrix}\boldsymbol{r}_{f}(\tau_{f}^{*})\\ \boldsymbol{v}_{f}(\tau_{f}^{*})\end{pmatrix}$ with corresponding final coast time $\tau_{f}^{*}$ iii. Calculate maximum constraint violation $\delta_{c}=\left\|\begin{pmatrix}\boldsymbol{r}(t)-\boldsymbol{r}_{f}(\tau_{f}^{*})\\ \boldsymbol{v}(t)-\boldsymbol{v}_{f}(\tau_{f}^{*})\end{pmatrix}\right\|_{\infty}$ iv. IF $\delta_{c}<\delta_{c,min}$ A. Update best values: $\delta_{c,min}=\delta_{c}$, $\tau_{s}=t$, $\tau_{f}=\tau_{f}^{*}$ (f) END FOR (g) IF $\delta_{c,min}<\delta$ i. Switch to local optimization routine with initial guess:
@@
First, the initial state and other fixed variables are set. The terminal boundary condition is defined to either be a target orbit or stable invariant manifold of an orbit. This provides an additional degree of freedom, described by the final coast time $\tau_{f}$, which characterizes the terminal point of the trajectory. In case of targeting an orbit, $\tau_{f}$ describes the time required to reach a fixed point on the orbit. When targeting the stable invariant manifold of a halo orbit $\tau_{f}$ corresponds to the coast time needed on the manifold to reach the insertion point on the halo orbit. After propagating the initial state (fixed) and initial costate (from the ACT method) according to the system dynamics, we determine the optimal terminal point for the given initial costate guess. To do so, for each state along the propagated trajectory, the nearest point on the target orbit or manifold is identified by iterating over the time of flight. This involves a nearest neighbor search based on the Euclidean distance of position and velocity [^1] Since $\boldsymbol{r}$ and $\boldsymbol{v}$ are normalized using natural units we use the combined $\left\|\cdot\right\|_{\infty}$-norm of both. In other cases these should be treated separately., which is conducted using a k-dimensional (k-d) tree [26]. This k-d tree only has to be generated once during the problem setup and is significantly more efficient than conventional search algorithms. After calculating the constraint violations relative to the closest target point for each state along the trajectory, the minimum constraint violation $\delta_{c,min}$ and the corresponding shooting time $\tau_{s}$ and final coast time $\tau_{f}$ are identified.
@@ 3.3 Generative Learning Framework
The framework used in this work to accelerate the global search problem of low-thrust trajectory optimization with generative machine learning is based on the Amortized Global Search (AmorGS) framework introduced by Li et al. [2]. The process, illustrated in Fig. 4, consists of three main steps. First, a large dataset of feasible or locally optimal solutions is generated using the indirect method combined with the global search strategy. These solutions correspond to a finite set of problems $\{\mathcal{P}_{\alpha}\}$, where the conditional parameter $\alpha$ is sampled from the a bounded set (e.g., $\alpha\in[\underline{\alpha},\overline{\alpha}]$). Next, the generated decision vectors $\overline{\boldsymbol{u}}$ and their associated $\alpha$ are used to train the generative machine learning model. In this paper a diffusion model is used for this purpose. A successful training of the model results in it learning how the solution space distribution varies with $\alpha$ and enables prediction of the structure for unseen values of the conditional variable. In the final step, the diffusion model generates initial guesses for a problem $\mathcal{P}_{\alpha}$, often corresponding to a conditional variable $\alpha$ that wasn’t encountered in the data generation and training phase. Ideally, these sampled indirect decision vectors $\overline{\boldsymbol{u}}$ serve as good approximations for unknown locally optimal solutions. Work to incorporate constraint satisfaction with diffusion models is nascent and not straightforward (see Li et al. [27] for some efforts on aerospace and robotic related problems). Hence the initial guesses generated by the diffusion model may not fully meet the desired feasibility tolerances. The initial guesses are therefore combined with the preliminary screening algorithm, which also calculates the shooting and final coast times $\tau_{s}$ and $\tau_{f}$ as described in Fig. 3. This has the additional benefit of the model not having to learn $\tau_{s}$ and $\tau_{f}$, therefore reducing the dimensionality of its input vector. After screening, the solutions can either be refined to local optimality using SNOPT or directly used as feasible solutions by setting a lower tolerance without further optimization.
The initial stages of generating data, training the model, and sampling from it are time-consuming and would not be necessary if data generation techniques were directly applied to the problem associated with the conditional parameter of interest. However, during preliminary mission design, the specific value of the conditional parameter might change frequently, requiring rapid generation of new solutions. As a result, it may be worth investing in the upfront effort to enable quick exploration and adaptation during early mission design phases or even in real-time during a mission to respond to unforeseen events.

### v2

@@ Abstract
The global search for optimal long time-duration, low-thrust spacecraft trajectories is a computationally expensive problem, that is characterized by clustering patterns in locally optimal solutions. During preliminary mission design, mission parameters have not been fully defined yet, necessitating that trajectory designers efficiently generate high-quality control solutions across a wide range of different scenarios. Generative machine learning models can be trained to learn how the solution structure varies with respect to an evolving mission parameter, thereby accelerating the global search for a large number of missions with varying parameters. In this work, state-of-the-art diffusion models are integrated with the indirect approach for trajectory optimization within a global search framework. The main difficulty with the indirect method lies in generating good initial guesses for the non-intuitive costate variables, which are crucial for solver convergence. By training a diffusion model to learn the structure of high-quality solutions in costate space, it can generate initial costate guesses for new mission parameters. This framework is tested on two low-thrust transfers of different complexity in the circular restricted three-body problem. By generating and analyzing a training data set, we develop mathematical relations and techniques to understand the complex structures in the costate domain of locally optimal solutions for these problems. A diffusion model is trained on this data and successfully predicts how the costate solution structure changes, based on the maximum spacecraft thrust magnitude. We warm-start a numerical solver with initial costates sampled from the diffusion model for problems with unseen thrust magnitudes and compare the number of solutions generated per minute to samples from a uniform distribution and from an adjoint control transformation. Results show that the diffusion model accelerates the global search process by one to two orders of magnitude, allowing for rapid generation of high-quality solutions.
@@ 1 Introduction
Low-thrust trajectory optimization is a global search problem with multiple objectives and constraints. Finding optimal solutions while minimizing fuel consumption and time of flight is a complex, computationally expensive and time-consuming endeavor tackled in the preliminary mission design phase. This task is further complicated by key mission parameters and objectives not being fixed yet. The flight-dynamics team therefore surveys a broad design space, generating hundreds to thousands of candidate trajectories to understand how performance varies across plausible parameter sets. For example, varying the maximum thrust magnitude can result in families of optimal trajectories with fundamentally different characteristics. Adjustments might also be required mid-mission, as exemplified by the BepiColombo spacecraft, operated by ESA and JAXA. In May 2024, reduced solar cell performance limited the thrusters’ operation to $90\%$ of maximum power [1]. Ideally, control solutions for such new constraints would be generated swiftly, either by ground engineers or autonomous on-board algorithms. We present a framework that uses generative machine learning to speed up the global search for trajectory optimization problems with varying mission parameters.
@@
Previous studies have demonstrated that locally optimal solutions to low-thrust trajectory optimization problems tend to form clusters [2, 3, 4, 5, 6]. Global search algorithms like monotonic basin hopping (MBH) aim to exploit this clustering to improve efficiency in generating solutions [7]. However, sampling distributions for MBH are typically chosen for their simplicity such as uniform, Gaussian, Cauchy, or Pareto distributions [8]. Their parameters often require manual tuning, although there has been work on adaptive self-tuning variants of MBH [9]. Li et al. [2] proposed an amortized global search framework that learns a conditional probability distribution supported on local basins of attraction for high-quality solutions using deep generative models. Their distribution was conditioned on the spacecraft’s maximum thrust, initially employing a conditional variational autoencoder (CVAE), later enhanced with a long short-term memory (LSTM), alongside a numerical solver to generate solutions for new thrust values. Beeson et al. [3] built on this work by mathematically defining the parameterized global search problem as the task of sampling a conditional probability distribution, which focuses on basins of attractions of high-quality solutions. Subsequent studies [10, 11] found that diffusion models (DM) outperformed the CVAE + LSTM approach for trajectory optimization. Diffusion probabilistic models, are state-of-the-art generative models, that have recently captured widespread interest, due to their strong performance in conditional image and video generation [12].
@@
The novel contribution of this work is the integration of a diffusion model framework with indirect optimal control for spacecraft trajectory optimization. While this builds on the work of Li et al. [2, 10], their research has focused solely on direct optimal control methods. Direct methods use transcription techniques to approximate dynamics and control, reformulating the optimal control problem as a nonlinear program. In contrast, indirect methods derive the first-order necessary conditions for optimality, where the additional costate variables are introduced. This results in a two-point boundary value problem. Although direct methods are now the standard in spacecraft trajectory optimization due to their simpler setup and robust convergence properties, especially with poor initial guesses, indirect methods provide a lower-dimensional solution space where each solution satisfies the first-order optimality conditions [13]. The main difficulty with the indirect method lies in generating good initial guesses for the non-intuitive costate variables, which are crucial for solver convergence. This challenge drives the use of generative learning in this context. By training a diffusion model to learn the structure of high-quality solutions in costate space, it can generate initial costate guesses for new conditional parameters. In general we use the expression conditional parameter to describe a physical parameter of the problem definition that, when varied, results in different families of trajectories.
@@
While the diffusion model facilitates the rapid generation of solutions for new conditional parameters, it requires a substantial amount of data for training. A key contribution of this work is the implementation of efficient methods to generate this training data. This is achieved by integrating the astrodynamics software pydylan [18] with adjoint control transformations (ACTs) and a modified version of the preliminary screening algorithm introduced by Russell [19]. The generative learning framework’s capabilities are demonstrated through two low-thrust transfers within the circular restricted three-body problem (CR3BP), with both transfers conditioned on the spacecraft’s maximum thrust. Training data is generated for fixed values of this parameter, and the diffusion model is then tested by generating solutions for a new thrust level not included in the training set. Results show that the diffusion model accelerates the global search process by one to two orders of magnitude, allowing for rapid generation of high-quality solutions.
@@ 2.1 Optimal Control Problem
The goal of optimal control theory is to determine a control signal which causes a process to satisfy a set of physical constraints, while maximizing or minimizing a performance metric [20]. This performance index is formulated as a cost or objective functional $J$, including a terminal cost $\phi$ and a running cost $\mathcal{L}$. A generic formulation of the problem is given by:
@@ 2.4 Circular Restricted Three Body Problem
Selecting a suitable dynamical model is vital to space mission design, as each strikes a different balance between realism and complexity. The circular restricted three-body problem (CR3BP) is commonly used during preliminary mission design for multi-body missions, as it simplifies the problem while utilizing dynamical structures for effective mission planning. In the CR3BP, a primary celestial body of mass $m_{1}$ and a secondary celestial body of mass $m_{2}$ ($m_{1}>m_{2}$) move in circular orbits around their common center of mass. The CR3BP describes the motion of a third body of negligible mass, that does not affect the orbits of the larger bodies [24]. Variables in the CR3BP are expressed using a system of normalized units, referred to as natural units (NU): the unit of mass is $m_{1}+m_{2}$, the distance unit (DU) is the constant distance between primary and secondary, and the time unit (TU) is chosen such that their orbital period around the barycenter is $2\pi$. The system’s only parameter is the mass ratio $\mu=\frac{m_{2}}{m_{1}+m_{2}}$. The third body’s motion is described in a rotating reference frame, where the primary and secondary are fixed along the $r_{1}$ -axis at $(-\mu,0,0)$ and $(1-\mu,0,0)$, with their barycenter at the origin. The $r_{3}$ -axis points in the direction of angular momentum, while the $r_{2}$-axis completes the right-handed system. The motion of the third body under the gravitational influence of $m_{1}$ and $m_{2}$ follows the dynamical system:
@@ 3.1 Adjoint Control Transformation
When using an indirect method, an initial guess for the costate variables is required to solve the boundary value problem. This is challenging due to the small convergence radius and lack of physical intuition for costates [26]. During data generation we employ a method referred to as adjoint control transformation (ACT) to generate initial costate guesses[27]. The ACT method maps a set of physical control variables to the non-physical costates: $\mathcal{M}:(\varphi,\dot{\varphi},\beta,\dot{\beta},S,\dot{S})\rightarrow(\boldsymbol{\lambda}_{r},\boldsymbol{\lambda}_{v})$, where $\varphi$ and $\beta$ represent the in-plane and out-of-plane angles of the thrust vector in a velocity-centered spacecraft coordinate frame. We evaluate this mapping at $t_{0}$ by uniformly sampling initial guesses for $\varphi_{0}$, $\beta_{0}$, $\dot{\varphi}_{0}$, $\dot{\beta}_{0}$, $S_{0}$ and $\dot{S}_{0}$ within problem-specific ranges, yielding reasonable initial guesses for $\boldsymbol{\lambda}_{r,0}$ and $\boldsymbol{\lambda}_{v,0}$. No mapping is needed for the mass costate at $t_{0}$, as we exploit the degree of freedom introduced by the factor $k$ by setting $\lambda_{m,0}=-1$. The transversality condition $\lambda_{m}(t_{f})=-k$ is satisfied, since $\dot{\lambda}_{m}\leq 0$.
[…]
@@ 3.2 Global Search Strategy
1. Set fixed values: $\boldsymbol{r}(0)=\boldsymbol{r}_{0}$, $\boldsymbol{v}(0)=\boldsymbol{v}_{0}$, $m(0)=m_{0}$, $\lambda_{m,0}=-1$, $\tau_{i}=0$ 2. Generate k-d tree for states on target orbit or manifold 3. DO WHILE ($j\leq j_{max}$ with $j\in\mathbb{N}$) (a) $j=j+1$ (b) Apply ACT method to obtain $\boldsymbol{\lambda}_{r,0}$ and $\boldsymbol{\lambda}_{v,0}$ (c) Propagate spacecraft states and costates according to Eqs. (6), (11) and (12) for $\tau_{s,max}$ and save $\boldsymbol{r}(t)$ and $\boldsymbol{v}(t)$ (d) $\delta_{c,min}=\infty$ (e) FOR $t=0$ to $t=\tau_{s,max}$ DO i. Query k-d tree to perform nearest neighbor search with $\begin{pmatrix}\boldsymbol{r}(t)\\ \boldsymbol{v}(t)\end{pmatrix}$ ii. k-d tree returns closest point on target orbit or manifold $\begin{pmatrix}\boldsymbol{r}_{f}(\tau_{f}^{*})\\ \boldsymbol{v}_{f}(\tau_{f}^{*})\end{pmatrix}$ with corresponding final coast time $\tau_{f}^{*}$ iii. Calculate maximum constraint violation $\delta_{c}=\left\|\begin{pmatrix}\boldsymbol{r}(t)-\boldsymbol{r}_{f}(\tau_{f}^{*})\\ \boldsymbol{v}(t)-\boldsymbol{v}_{f}(\tau_{f}^{*})\end{pmatrix}\right\|_{\infty}$ iv. IF $\delta_{c}<\delta_{c,min}$ A. Update best values: $\delta_{c,min}=\delta_{c}$, $\tau_{s}=t$, $\tau_{f}=\tau_{f}^{*}$ (f) END FOR (g) IF $\delta_{c,min}<\delta$ i. Switch to local optimization routine with initial guess:
@@
First, the initial state and other fixed variables are set. The terminal boundary condition is defined to either be a target orbit or stable invariant manifold of an orbit. This provides an additional degree of freedom, described by the final coast time $\tau_{f}$, which characterizes the terminal point of the trajectory. In the case of targeting an orbit, $\tau_{f}$ describes the time required to reach a fixed point on the orbit. When targeting the stable invariant manifold of a halo orbit, $\tau_{f}$ corresponds to the coast time needed to reach the insertion point. After propagating the initial state (fixed) and initial costate (from the ACT method) according to the system dynamics, we determine the optimal terminal point for the given initial costate guess. To do so, for each state along the propagated trajectory, the nearest point on the target orbit or manifold is identified by iterating over the time of flight. This involves a nearest neighbor search based on the Euclidean distance of position and velocity [^1] Since $\boldsymbol{r}$ and $\boldsymbol{v}$ are normalized using natural units we use the combined $\left\|\cdot\right\|_{\infty}$-norm of both. In other cases these should be treated separately., which is conducted using a k-dimensional (k-d) tree [29]. This k-d tree only has to be generated once during the problem setup and partitions the states of the target orbit or manifold in a tree structure. It allows the nearest neighbor search for a given trajectory to be conducted more efficiently than conventional search algorithms, by moving down the tree structure recursively. After calculating the constraint violations relative to the closest target point for each state along the trajectory, the minimum constraint violation $\delta_{c,min}$ and the corresponding shooting time $\tau_{s}$ and final coast time $\tau_{f}$ are identified.
@@ 3.3 Generative Learning Framework
The framework used in this work to accelerate the global search problem of low-thrust trajectory optimization with generative machine learning is based on the Amortized Global Search (AmorGS) framework introduced by Li et al. [2]. The process, illustrated in Fig. 4, consists of three main steps. First, a large dataset of feasible or locally optimal solutions is generated using the indirect method combined with the global search strategy. These solutions correspond to a finite set of problems $\{\mathcal{P}_{\alpha}\}$, where the conditional parameter $\alpha$ is sampled from a bounded set (e.g., $\alpha\in[\underline{\alpha},\overline{\alpha}]$). Next, the generated decision vectors $\overline{\boldsymbol{u}}$ and their associated $\alpha$ are used to train the generative machine learning model, as shown at the top of Fig. 4. In this paper a diffusion model is used for this purpose. A successful training of the model results in it learning how the solution space distribution (visualized in Fig. 5) varies with $\alpha$ and enables prediction of the structure for unseen values of the conditional variable. In the final step, labeled Model Testing in Fig. 4, the diffusion model generates initial guesses for a problem $\mathcal{P}_{\alpha}$, often corresponding to a conditional variable $\alpha$ that wasn’t encountered in the data generation and training phase. Ideally, these sampled indirect decision vectors $\overline{\boldsymbol{u}}$ serve as good approximations for unknown locally optimal solutions. Work to incorporate constraint satisfaction with diffusion models is nascent and not straightforward (see Li et al. [30] for some efforts on aerospace and robotic related problems). Hence, the initial guesses generated by the diffusion model may not fully meet the desired feasibility tolerances. The initial guesses are therefore combined with the preliminary screening algorithm, which also calculates the shooting and final coast times $\tau_{s}$ and $\tau_{f}$ as described in Fig. 3. This has the additional benefit of the model not having to learn $\tau_{s}$ and $\tau_{f}$, therefore reducing the dimensionality of its input vector. After screening, the solutions can either be refined to local optimality using SNOPT (upper path in Fig. 4) or directly used as feasible solutions without further optimization by setting the global search tolerance to the desired feasibility tolerance (lower path in Fig. 4).
The initial stages of generating data (5,000-10,000 CPU-hours), training the model (3-4 GPU-hours), and sampling from it (5-6 GPU-minutes) are time-consuming and would not be necessary if we directly solve the target problem. However, during preliminary mission design, the specific value of the conditional parameter is not fixed yet, requiring rapid generation of solutions across wide parameter ranges. As a result, it may be worth investing in the upfront effort to enable quick exploration and adaptation during early mission design phases or even in real-time during a mission to respond to unforeseen events.